Venture capital. Henry invested $12,000 in a new
restaurant. When the restaurant was sold two years
later, he received $27,000. Find his average annual
return by solving the equation 12,000(1 + r)^2=
27,000.

Answer by tutorcecilia(2152)

I am assumeing that "r" is the average annual revenue, so I will solve for "r":
12,000(1 + r)^2=27,000
12,000/12,000(1+r)^2=27,000/12,000 [Isolate the r-variable]
(1 + r)^2=2.25
(sqrt(1+r)^2=sqrt(2.25)
1+r=1.5
r=1.5-1
r=0.5
Check by plugging 0.5 back into the original equation and solve.